OpenStudy (anonymous):

What is an equation of the line in slope-intercept form? m=3 and the y-intercept is (0, 4)

OpenStudy (anonymous):

Choices: y=4x+3 y=-3x+4 y=4x-3 y=3x+4

OpenStudy (isaiah.feynman):

Easy peasy. y=mx+b. Use that form to find the answer.

OpenStudy (anonymous):

slope = 4 y =mx+c put the point (0,4) in the eq and get your answer

jimthompson5910 (jim_thompson5910):

as the previous answers are saying, use y = mx+b where m = slope b = y-intercept

OpenStudy (anonymous):

I used y=mx+b. I ended up getting 3x=-3. I don't know how.

jimthompson5910 (jim_thompson5910):

the given slope is m = 3 so we go from y = mx+b to y = 3x+b after we replace m with 3

jimthompson5910 (jim_thompson5910):

the given y-intercept is (0,4) or just 4 so b = 4 and we replace b with 4 to go from y = 3x+b to y = 3x+4

OpenStudy (anonymous):

y=mx+c where c is the y intercept... just place the values..

OpenStudy (isaiah.feynman):

Looks like you dunno what the term in that equation represent.